OpenStudy (anonymous):

y=(x-1)^2+2 graph each function

OpenStudy (lucaz):

this is in the form y=(x-h)^2+k (h,k) is the vertex

OpenStudy (lucaz):

since the coefficient of (x-h)^2 is positive this parabola opens upward

OpenStudy (anonymous):

oo i see

OpenStudy (anonymous):

y = (x-1)(x-1) +2 y = x^2-2x+1+2 y = x^2-2x+3 let y= some number {-1, 0, etc}, and solve for x

OpenStudy (anonymous):

Because it is a parabola, x=1 will be mirrored over the y-axis in quad 2

OpenStudy (anonymous):

as well as all other x's

OpenStudy (anonymous):

ok